It was proposed that a study be conducted on the number of start- ups and annual number of business bankruptcies. Data collected was six year period. Develop regression equation and forecast the bankruptcy numbers given the start ups for the forthcoming year to be 60. Interpret the significance of predictor through R Square


Business Bankruptcies Start ups

(in 1000)

34.3 58.1

35 55.4

38.5 57

40.1 58.5

35.5 57.4

37.9 58


1
Expert's answer
2021-02-24T07:28:27-0500

Answer


Working





Regression Line


The mean is given by

"\bar{x}"= "1 \over n""\u03a3x"


bankruptcies

"\bar{x}" = "344.4 \over 6" = 57.4


Startups

"\bar{y}"= "221.3 \over 6" =36.8833

The slope is given by


β1= "Σ (x-\bar{x}) (y-\bar{y}) \over Σ (x-\bar{x})^2"


β1= "5.46 \over 6.22"= 0.8778


The intercept is given by


 β0= "\bar{y}"β1"\bar{x}"


β0= 36.8833-57.4 x 0.8778 = -13.5032


The regression line is given by

y=-13.5032+ 1.17725x



When start ups = 60, then predicted bankruptcies are

y = -13.5032 + 0.8778 x 60

y = 39.1656



Error sum of square is given by


ESS = "\u03a3 (Residual)^2" = 21.3355


Total sum of square is given by


TSS= "Σ (y-\bar{y})^2" = 26.1283


The r-square is given by


R2=1-"ESS \over TSS"



R2=1-"21.3355 \over 26.1283"


R2= 0.1834


Interpretation

It shows that around 18.34% variation in the bankruptcies is explained by the number of start-ups.